# Bath-induced interactions and transient dynamics in open quantum systems at strong coupling: Effective Hamiltonian approach

### v1

@@ II.4 Eigenenergies of the effective Hamiltonian
We have shown that the effect of the bath becomes embedded into the effective Hamiltonian $\hat{H}^{\rm eff}_{S}$ in Eq. (9). The thermal energy from the bath can be used to induce transitions between the energy levels of the effective Hamiltonian. A simulation of the dynamics considering the reaction-coordinate Hamiltonian, Eq. (II.2), contains many of these transitions; however, whenever $\Omega\gtrsim\lambda,T$, we expect the effective Hamiltonian to capture the most important transitions. As an example of the power of the effective approach, we show below that we can infer the bath-induced interaction energy between spins from the dynamical signals. While this is difficult in general, including in the reaction-coordinate treatment, the effective Hamiltonian introduces a transparent way to account for inter-spin interactions induced by the bath.
@@
At small $\lambda$, the four eigenvalues can be associated with the energies of the free spins, with the two spins pointing down with eigenenergy $\omega_{-+}$, two single-excitation states with energies $\omega_{--}$ and $\omega_{+-}$, and the doubly excited state with energy $\omega_{++}$, where
$\omega_{-+}$ $\xrightarrow{\lambda\to 0}-(\Delta_{2}+\Delta_{1}),$
$\omega_{--}$ $\xrightarrow{\lambda\to 0}-\Delta_{1}+\Delta_{2},$
$\omega_{+-}$ $\xrightarrow{\lambda\to 0}\Delta_{1}-\Delta_{2},$
$\omega_{++}$ $\xrightarrow{\lambda\to 0}(\Delta_{2}+\Delta_{1}).$ (19)
@@
III Results
We assess the effective Hamiltonian method in its ability to produce the correct dynamics and steady-state results for a broad range of system-bath coupling strengths. The effective Hamiltonian method has been benchmarked so far only for equilibrium Anto-Sztrikacs et al. (2023a, b) and nonequilibrium steady state problems Anto-Sztrikacs et al. (2023a), with a focus on impurity models. Lattice models were examined in Ref. Min et al. (2023). Here, we focus on system-bath configurations including several spins and test the capability of the method to capture the polarization behavior throughout the entire range of coupling in both the transient and steady-state regimes. Our focus is in particular on capturing and quantifying bath-induced interactions through the effective Hamiltonian method.
Sec. III.1 and Sec. III.2 are devoted to equilibrium properties, the magnetic polarization and correlations, respectively. We analyze the out-of-equilibrium dynamics for the cases $N=1$ and $N=2$ in Sec. III.3 and Sec. III.4, respectively. In Sec. III.5 we analyze the dynamics of correlations between the two spins for the $N=2$ case.
Based on our results, we classify the following coupling regimes: The ultraweak coupling regime corresponds to $\lambda/\Delta\ll 1$; the weak to intermediate coupling region refers to $\lambda/\Delta\lesssim 3$. In the strong coupling regime, extending $3\lesssim\lambda/\Delta\lesssim 10$, the dynamics and the equilibrium state strongly deviate from the conventional Born-Markov perturbative result. The ultrastrong coupling limit corresponds to $\lambda/\Delta\gtrsim 10$. In establishing these regimes we have assumed that in multi-spin scenarios, the spin splittings are similar, i.e., $\Delta_{i}\approx\Delta$.
III.1 Polarization at thermal equilibrium
Figure 3: Average magnetization at thermal equilibrium for the $N=1$ (grey), $N=2$ (blue) and $N=3$ (green) cases as a function of coupling energy $\lambda$. Circles denote the calculation for the reaction-coordinate Hamiltonian $\hat{H}_{S}^{\rm RC}$ ($M=50$) and squares the equivalent calculation using the effective Hamiltonian $\hat{H}_{S}^{\rm eff}$. The parameters for the calculation are $\Omega=8\Delta_{1}$, $\Delta_{2}=0.9\Delta_{1}$ and $T=\Delta_{1}$. HEOM calculations are shown with parameters $N_{c}=5$ and $N_{k}=10$ (see Appendix C).
@@ III.1 Polarization at thermal equilibrium
with $Z^{\rm eff}=\mathrm{Tr}[e^{-\beta\hat{H}_{S}^{\rm eff}}]$. The expectation value of the polarization for each of the spins composing the $N=2$ system follows from
@@
where $\bullet$ denotes either the reaction-coordinate expectation value or its effective Hamiltonian counterpart.
In Fig. 3, we present the average magnetization, $\sum_{i}\langle\hat{\sigma}^{z}_{i}\rangle^{\rm RC}/N$ and $\sum_{i}\langle\hat{\sigma}^{z}_{i}\rangle^{\rm eff}/N$ at thermal equilibrium as a function of the coupling energy parameter $\lambda$. Together with these results, we complement the calculation with the expectation values obtained from HEOM simulations for certain values of $\lambda$. We make the following observations:
@@
(iii) The effective Hamiltonian method becomes more accurate as the number of spins $N$ is increased, in the sense that its results more closely approach those from the RC method and HEOM. The polarization is already well-approximated, and the behavior is effectively captured in the two-spin model through our effective Hamiltonian treatment. Interestingly, increasing the system complexity by adding more spins translates to an increasingly better approximation from the effective Hamiltonian.
@@ III.2 Correlations at thermal equilibrium
Figure 4: Quantum mutual information at thermal equilibrium between each of the spins composing the two-spin system as a function of coupling energy $\lambda$. Circles denote the calculation for the reaction-coordinate Hamiltonian $\hat{H}_{S}^{\rm RC}$ ($M=50$) and squares the equivalent calculation using the effective Hamiltonian $\hat{H}_{S}^{\rm eff}$. The parameters for the calculation are $\Omega=8\Delta_{1}$ and $\Delta_{2}=0.9\Delta_{1}$. Three values of temperature $T=1/\beta$ are shown from left to right: $T=\Delta_{1},2\Delta_{1},4\Delta_{1}$. HEOM calculations are shown with parameters $N_{c}=5$ and $N_{k}=10$ (see Appendix C).
To address this objective, we consider again the equilibrium state as described in Sec. III.1. We intend to understand if the effective Hamiltonian provides a good approximation of this state and hence the correlations among each subsystem composing the two-spin model.
@@
III.3 Dynamics $N=1$
@@
We now turn to the dynamics of a single spin coupled to a bath starting from an out-of-equilibrium state. We recall that Eq. (II.3) dictates the dynamics of the reduced state of a spin system resulting from the interaction with the bath.
@@
In Fig. 5 we show the dynamics of the polarization as produced by the reaction coordinate method (blue) compared to the ones obtained from the effective Hamiltonian (red) and the weak-coupling Redfield equation limit (gray) for two values of coupling energy, $\lambda=\Delta_{1},2.5\Delta_{1}$, $T=\Delta_{1}$ and $\Omega=8\Delta_{1}$. It is observed that the transient dynamics are not accurately captured by our effective Hamiltonian model. For the case of $N=1$, the observed oscillations in the spin polarization are due to non-Markovian signatures in the relaxation dynamics that are not included in our effective Hamiltonian. We thus find that for the $N=1$ case, while we can understand the equilibrium (long-time) properties from the effective Hamiltonian (Sec. III.1), the transient dynamics require further refinement of our effective treatment. Specifically, the complete truncation of the RC manifold down to its ground state erases non-Markovian signatures in the relaxation dynamics. Maintaining more levels in the RC manifold should recover those features.
@@ III.4 Dynamics $N=2$
In Fig. 6, Fig. 7 and Fig. 8, we display the dynamics of $\langle\hat{\sigma}^{z}_{i}\rangle(t)$ for the values of $\lambda=\Delta_{1}$, $\lambda=2.5\Delta_{1}$ and $\lambda=5\Delta_{1}$; respectively. For these calculations, we chose $\Omega=8\Delta_{1}$ and $\Delta_{2}=0.9\Delta_{1}$, for a particular value of temperature $T=\Delta_{1}$. The dynamics is characterized by oscillations until the stationary value for the magnetization is reached. We also display the dynamics of the $N=1$ case with the specific value of $\Delta_{i}$ for the $i$-th spin. Trivially, synchronization does not occur in this case due to the absence of the other spin degree of freedom. Furthermore, in Fig. 6, we display the dynamics using the weak-coupling approach for the $N=2$ model, which neither posses the aforementioned oscillations in the spin polarization due to spin-bath effects, nor the synchronization effect observed due to strong-coupling and bath-mediated interactions. We recall that the weak-coupling calculation is carried out with the Brownian spectral density function in Eq. (15), as opposed to the Ohmic spectrum in Eq. (16).
We find that the transient dynamics is well-approximated by our effective Hamiltonian treatment in the regime of weak to intermediate coupling. Most importantly, Fig. 6 and Fig. 7 demonstrate that the effective Hamiltonian treatment captures both the frequency and magnitude of the bath-induced synchronization dynamics. At stronger coupling, Fig. 8 shows that while the effective method properly captures the period of synchronization, it overestimates the magnitude of oscillations, or in other words, it underestimates the dissipative dynamics.
@@
In arriving at Eq. (30), we recall that, as $\lambda\to 0$, each of the four eigenvalues of the effective Hamiltonian correspond to the eigenstates $\ket{\downarrow\downarrow}$, $\ket{\uparrow\downarrow}$, $\ket{\downarrow\uparrow}$ and $\ket{\uparrow\uparrow}$ (see Eq. (II.4) and Fig. 2). These eigenvalues change as $\lambda$ increases due to the effect of the bath. Given that $\hat{S}=\hat{\sigma}^{x}_{1}+\hat{\sigma}^{x}_{2}$, the bath can induce spin-flipping interactions on both spins but it does not admix spin occupation sectors. In light of this, the choice of initial condition becomes important. With our preparation, $\hat{\rho}(t=0)=\ket{\uparrow\downarrow}\bra{\uparrow\downarrow}$ |, the relevant mode presented in the dynamics will be the one given by the difference between the states corresponding to $\ket{\uparrow\downarrow}$ and $\ket{\downarrow\uparrow}$, in such a way that the frequency of oscillations must be related to $\delta\omega_{\rm eff}$ in Eq. (30).
@@
We have found that the effective Hamiltonian model, while capable of approximating the multi-spin synchronized dynamics at weak-to-moderate couplings (Fig. 6-7) for special (experimentally-relevant) initial conditions, fails to replicate the exact dynamics at stronger coupling. However, while the amplitude of oscillations is not appropriately captured by the effective Hamiltonian, the periodicity in oscillations is correct, and it gives an analytical understanding of relevant energy scales. From Eq. (31), we observe that these oscillations are characterized by a frequency which is temperature independent.
@@
III.5 Dynamics of correlations $N=2$
In Sec. III.2, we have observed that the effect of the bath onto the two-spin system $N=2$ is to generate correlations at thermal equilibrium. It is of interest to consider the generation of these correlations and, in particular, to analyze the nature of these correlations.
@@ III.5 Dynamics of correlations $N=2$
Figure 10: Dynamics of the QMI (top) and entanglement negativity (bootm) as a function of time for the two-spin system ($\Delta_{2}=0.9\Delta_{1}$) obtained from the (blue) reaction-coordinate mapping ($M=25$) and the (red) effective Hamiltonian treatment for $T=\Delta_{1}$ and $\lambda=\Delta_{1}$. Calculations are shown for $\lambda=\Delta_{1}$. The dynamics is computed from the Redfield QME in Eq. (II.3). The residual coupling parameter is $\gamma=0.05$, [see Eq. (16)].
@@
Most importantly, we see that the dynamics reveal that there are correlations developing in the transient states from the QMI [Fig. 10(top)], which prevail even in the limit of long times (see Fig. 4). Interestingly, from the entanglement negativity [Fig. 10(bottom)], it is observed that some of these correlations as a function of time are non-classical in nature, meaning that they stem from the non-separability of the reduced state between spin $A$ and spin $B$. However, from the computation of the entanglement negativity of the equilibrium states (an analogous calculation of the entanglement negativity as in Fig. 4) we have obtained that in the equilibrium states the entanglement negativity is zero. This immediately implies that all the correlations in Fig. 4 are classical correlations. There are non-classical correlations in the transient states, as observed in Fig. 10(bottom), but they are washed away in the limit of long times. The presence of this non-classical correlations may be relevant to the engineering of entangled states for quantum computation. We highlight that these intriguing observations may not hold at lower temperatures, a question that we leave for future studies.
@@
IV Assessment of the effective Hamiltonian method
We summarize the regime of applicability of the effective Hamiltonian theory. The effective Hamiltonian method prepares a system Hamiltonian that incorporates strong coupling effects through, e.g., renormalization of energy parameters and the generation of bath-induced interaction terms. As such, by inspecting the effective Hamiltonian one can gain a direct understanding of expected strong coupling signatures. On the other hand, the coupling of the effective Hamiltonian to the residual bath is assumed weak and it is handled at the level of the Born-Markov approximation. Therefore, non-Markovian relaxation effects in the effective Hamiltonian dynamics are not taken into account. Based on this understanding, we summarize our assessment of the effective Hamiltonian theory:
(i) The theory properly captures the steady state behavior of spins in a boson bath, from the ultraweak to the ultrastrong coupling limits. Significantly, it becomes more accurate when increasing the number of spins in the system.

### v2

@@ II.4 Eigenenergies of the effective Hamiltonian
We have shown that part of the effect of the bath becomes embedded into the effective Hamiltonian $\hat{H}^{\rm eff}_{S}$ in Eq. (9). The thermal energy from the bath can be used to induce transitions between the energy levels of the effective Hamiltonian. Simulations of the dynamics considering the reaction-coordinate Hamiltonian, Eq. (II.2), contain many of these transitions; however, whenever $\Omega\gtrsim\lambda,T$, we expect the effective Hamiltonian to capture the most important transitions. The intuition behind this assumption rests on the fact that the truncation performed on the energy manifold of the reaction coordinate degree of freedom in the polaron frame introduces an error [see Eq. (II.2)]. However, we expect this error to remain small whenever the population of the excited states in the reaction-coordinate manifold is small. The physical reasoning behind this assumption is that the reaction-coordinate levels are sufficiently spaced out whenever $\Omega\gtrsim\lambda$ and the population of the highly-excited levels of the reaction coordinate will be small. The second assumption regarding energy scales is that thermal fluctuations will not lead to large populations of the highly-excited states of the reaction coordinate in the polaron frame, that is, $\Omega\gtrsim T$.
Interestingly, it has been demonstrated that the $\Omega\gtrsim\lambda$ condition can be relaxed for certain lattice models such that the effective model leads to an excellent approximation even in the case of arbitrarily large $\lambda$ Min et al. (2023). However, for the purpose of evaluating the dynamics in this work, we restrict ourselves to the parameter regime in the energy space where the effective model is straightforwardly justified.
As an example of the power of the effective approach, we show below that we can infer the bath-induced interaction energy between spins from the dynamical signals. While this is difficult in general to achieve with numerical tools (including in the reaction-coordinate treatment), the effective Hamiltonian introduces a transparent way to account for inter-spin interactions induced by the bath.
@@
At small $\lambda$, the four eigenvalues can be associated with the energies of the free spins, with the two spins pointing down with eigenenergy $\omega_{-,+}$, two single-excitation states with energies $\omega_{-,-}$ and $\omega_{+,-}$, and the doubly excited state with energy $\omega_{+,+}$, where
$\omega_{-,+}$ $\xrightarrow{\lambda\to 0}-(\Delta_{2}+\Delta_{1}),$
$\omega_{-,-}$ $\xrightarrow{\lambda\to 0}-\Delta_{1}+\Delta_{2},$
$\omega_{+,-}$ $\xrightarrow{\lambda\to 0}\Delta_{1}-\Delta_{2},$
$\omega_{+,+}$ $\xrightarrow{\lambda\to 0}(\Delta_{2}+\Delta_{1}).$ (19)
@@
III Results: Thermal equilibrium
In this Section, we assess the effective Hamiltonian method in its ability to produce the correct steady-state results for a broad range of system-bath coupling strengths. The effective Hamiltonian method has been benchmarked so far only for equilibrium Anto-Sztrikacs et al. (2023a, b) and nonequilibrium steady state problems Anto-Sztrikacs et al. (2023a), with a focus on impurity models. Lattice models have been examined in Ref. Min et al. (2023). Here, we focus on system-bath configurations including few spins and test the capability of the method to capture the polarization behavior throughout the entire range of coupling in the steady-state regime (transients are explored in Sec. IV).
Figure 3: Average magnetization at thermal equilibrium for the $N=1$ (grey), $N=2$ (blue) and $N=3$ (green) cases as a function of coupling energy $\lambda$. Circles denote the calculation for the reaction-coordinate Hamiltonian $\hat{H}_{S}^{\rm RC}$ ($M=50$) and squares the corresponding calculation using the effective Hamiltonian $\hat{H}_{S}^{\rm eff}$. The parameters for the calculation are $\Omega=8\Delta_{1}$, $\Delta_{2}=0.9\Delta_{1}$ and $T=\Delta_{1}$. HEOM calculations are shown with parameters $N_{c}=5$ and $N_{k}=10$ (see Appendix C).
@@ III Results: Thermal equilibrium
with $Z^{\rm eff}=\mathrm{Tr}[e^{-\beta\hat{H}_{S}^{\rm eff}}]$.
III.1 Polarization
The expectation value of the polarization for each of the spins composing the $N=2$ system follows from
@@ III.1 Polarization
where $\bullet$ denotes either the reaction-coordinate expectation value or its effective Hamiltonian counterpart. In Fig. 3, we present the average magnetization, $\sum_{i}\langle\hat{\sigma}^{z}_{i}\rangle^{\rm RC}/N$ and $\sum_{i}\langle\hat{\sigma}^{z}_{i}\rangle^{\rm eff}/N$ at thermal equilibrium as a function of the coupling energy parameter $\lambda$. Together with these results, we complement the calculation with the expectation values obtained from HEOM simulations for certain values of $\lambda$. We make the following observations:
@@
(iii) It appears that the effective Hamiltonian method becomes more accurate with respect to the RC and HEOM approaches moving from $N=1$ to $N=3$. The polarization is already well-approximated, and the behavior is effectively captured in the two-spin model through our effective Hamiltonian treatment.
@@ III.2 Correlations
Figure 4: Quantum mutual information at thermal equilibrium between each of the spins composing the two-spin system as a function of coupling energy $\lambda$. Circles denote the calculation for the reaction-coordinate Hamiltonian $\hat{H}_{S}^{\rm RC}$ ($M=50$) and squares the corresponding calculation using the effective Hamiltonian $\hat{H}_{S}^{\rm eff}$. The parameters for the calculation are $\Omega=8\Delta_{1}$ and $\Delta_{2}=0.9\Delta_{1}$. Three values of temperature $T=1/\beta$ are shown from left to right: $T=\Delta_{1},2\Delta_{1},4\Delta_{1}$. HEOM calculations are shown with parameters $N_{c}=5$ and $N_{k}=10$ (see Appendix C).
To address this objective, we consider again the equilibrium state as described in Eq. (20). We intend to understand if the effective Hamiltonian provides a good approximation of this state and hence the correlations among each subsystem composing the two-spin model.
@@
IV Results: Dynamics
We analyze the out-of-equilibrium dynamics for the cases $N=1$ and $N=2$ in Sec. IV.1 and Sec. IV.2, respectively. For two spins, we investigate their synchronization in Sec. IV.2.1 and the dynamics of correlations in Sec. IV.2.2. With respect to the dynamics from the HEOM method described in Sec. II.3, we demonstrate in Appendix C that tier (i) and tier (ii) methods, HEOM and RC simulations, respectively, provide the same results. For this reason, for dynamical features we only display results from the cheaper approximation tier (ii). Our focus is in particular on capturing and quantifying bath-induced interactions through the effective Hamiltonian method.
IV.1 Single spin: Relaxation to equilibrium
@@
We focus here on the dynamics of a single spin coupled to a bath starting from an out-of-equilibrium state. We recall that Eq. (II.3) dictates the dynamics of the reduced state of a spin system resulting from the interaction with the bath.
@@
In Fig. 5 we show the dynamics of the polarization as produced by the reaction coordinate method (blue) compared to the ones obtained from the effective Hamiltonian (red) and the weak-coupling Redfield equation limit (gray) for two values of coupling energy, $\lambda=\Delta_{1},2.5\Delta_{1}$, $T=\Delta_{1}$ and $\Omega=8\Delta_{1}$. It is observed that the transient dynamics are not accurately captured by our effective Hamiltonian model. For the case of $N=1$, the observed oscillations in the spin polarization are due to non-Markovian signatures in the relaxation dynamics that are not included in our effective Hamiltonian. We thus find that for the $N=1$ case, while we can understand the equilibrium (long-time) properties from the effective Hamiltonian (Sec. III.1), the transient dynamics require further refinement of our effective treatment. Specifically, the complete truncation of the RC manifold down to its ground state erases non-Markovian signatures in the relaxation dynamics. Maintaining more levels in the RC manifold could recover those features.
@@ IV.2 Two spins: Synchronization and Correlations
Our objective is to evaluate the dynamics generated by the effective Hamiltonian and compare them with exact results. In Appendix C, we show that the dynamics generated from the full reaction-coordinate Hamiltonian coincide with those generated with HEOM (tiers (i) and (ii) from Sec. II.3). For this reason, we shall consider the full reaction-coordinate Hamiltonian as exact results.
IV.2.1 Synchronization behavior and its temperature dependence
In Fig. 6, Fig. 7 and Fig. 8, we display the dynamics of $\langle\hat{\sigma}^{z}_{i}\rangle(t)$ for the values of $\lambda=\Delta_{1}$, $\lambda=2.5\Delta_{1}$ and $\lambda=5\Delta_{1}$; respectively. For these calculations, we chose $\Omega=8\Delta_{1}$ and $\Delta_{2}=0.9\Delta_{1}$, for a particular value of temperature $T=\Delta_{1}$. The blue curves depict the dynamics generated by the reaction-coordinate Hamiltonian, i.e., the exact results; while the red curves display the dynamics generated with the effective Hamiltonian model.
The dynamics is characterized by oscillations until the stationary value for the magnetization is reached. We also display the dynamics of the $N=1$ case with the specific value of $\Delta_{i}$ for the $i$-th spin. Trivially, synchronization does not occur in this case due to the absence of an interaction with any other spin degree of freedom This can be observed in the solid-gray curves in Fig. 6, Fig. 7 and Fig. 8. Furthermore, in Fig. 6, we display the dynamics using the weak-coupling approach for the $N=2$ model with dashed-gray curves, which neither posses the aforementioned oscillations in the spin polarization due to spin-bath effects, nor the synchronization effect observed due to strong-coupling and bath-mediated interactions Lam . As a reminder, the weak-coupling calculation is carried out with the Brownian spectral density function in Eq. (15), as opposed to the Ohmic spectrum achieved after the RC mapping, Eq. (16).
We find that the transient dynamics is well-approximated by our effective Hamiltonian treatment in the regime of weak to intermediate coupling. Most importantly, Fig. 6 and Fig. 7 demonstrate that the effective Hamiltonian treatment captures both the frequency and magnitude of the bath-induced synchronization dynamics. At stronger coupling, Fig. 8 shows that while the effective method properly captures the period of synchronization, it overestimates the magnitude of oscillations, or in other words, it underestimates the dissipative dynamics. Synchronization, in this case, becomes apparent from the common frequency of the magnetization oscillations in time for both spins. Such synchronized dynamics arise from the interaction induced by the bath and not due to internal degrees of freedom.
@@ IV.2.1 Synchronization behavior and its temperature dependence
In arriving at Eq. (30), we recall that, as $\lambda\to 0$, each of the four eigenvalues of the effective Hamiltonian correspond to the eigenstates $\ket{\downarrow\downarrow}$, $\ket{\uparrow\downarrow}$, $\ket{\downarrow\uparrow}$ and $\ket{\uparrow\uparrow}$ (see Eq. (II.4) and Fig. 2). These eigenvalues change as $\lambda$ increases due to the effect of the bath. Following $\hat{H}^{\rm{eff}}_{S}$ from Eq. (9) and $\hat{S}=\hat{\sigma}^{x}_{1}+\hat{\sigma}^{x}_{2}$, after ignoring constant terms to the Hamiltonian we have
$\hat{H}^{\rm{eff}}_{S}$ $=e^{-\frac{2\lambda^{2}}{\Omega^{2}}}\left(\Delta_{1}\hat{\sigma}^{z}_{1}+\Delta_{2}\hat{\sigma}^{z}_{2}\right)-\frac{2\lambda^{2}}{\Omega}\hat{\sigma}^{x}_{1}\hat{\sigma}^{x}_{2},$ (32)
where the second term corresponds to the interaction induced by the common reservoir. This Hamiltonian is $\mathbb{Z}_{2}$ symmetric, such that the operator $\hat{O}=\prod_{i=1}^{N}\hat{\sigma}^{z}_{i}$ commutes with the Hamiltonian. This implies that the bath effective interaction energy can induce spin-flipping interactions on both spins but it does not admix spin inversion sectors. In light of this, the choice of initial condition becomes important. With our preparation, $\hat{\rho}(t=0)=\ket{\uparrow\downarrow}\bra{\uparrow\downarrow}$ |, the relevant mode presented in the dynamics will be the one given by the difference between the states corresponding to $\ket{\uparrow\downarrow}$ and $\ket{\downarrow\uparrow}$, both of which belong to the same spin inversion subsector. In such a way that the frequency of oscillations must be related to $\delta\omega_{\rm eff}$ in Eq. (30).
@@
We have found that the effective Hamiltonian model, while capable of approximating the multi-spin synchronized dynamics at weak-to-moderate couplings (Fig. 6-7) for special (experimentally-relevant) initial conditions, fails to replicate the exact dynamics at stronger coupling. However, while the amplitude of oscillations is not appropriately captured by the effective Hamiltonian, the periodicity in oscillations is correct, and it gives an analytical understanding of relevant energy scales.
From Eq. (31), we observe that these oscillations are characterized by a frequency which is temperature independent.
@@
IV.2.2 Dynamics of correlations
In Sec. III.2, we have observed that the effect of the bath onto the two-spin system $N=2$ is to generate correlations at thermal equilibrium. It is of interest to consider the dynamics of these correlations and, in particular, to analyze the nature of these correlations.
@@ IV.2.2 Dynamics of correlations
Figure 10: Dynamics of the QMI (top) and entanglement negativity (bottom) as a function of time for the two-spin system ($\Delta_{2}=0.9\Delta_{1}$) obtained from the (blue) reaction-coordinate mapping ($M=25$) and the (red) effective Hamiltonian treatment for $T=\Delta_{1}$ and $\lambda=\Delta_{1}$. Calculations are shown for $\lambda=\Delta_{1}$. The dynamics is computed from the Redfield QME in Eq. (II.3). The residual coupling parameter is $\gamma=0.05$, [see Eq. (16)].
@@
Most importantly, we see that the dynamics reveal that there are correlations developing in the transient states from the QMI [Fig. 10(top)], which prevail even in the limit of long times (see Fig. 4). Interestingly, from the entanglement negativity [Fig. 10(bottom)], it is observed that some of these correlations as a function of time are non-classical in nature, meaning that they stem from the non-separability of the reduced state between spin $A$ and spin $B$. However, from the computation of the entanglement negativity of the equilibrium states (an analogous calculation of the entanglement negativity as in Fig. 4) we have obtained that in the equilibrium states the entanglement negativity is zero. This immediately implies that all the correlations in Fig. 4 are classical correlations. There are non-classical correlations in the transient states, as observed in Fig. 10(bottom), but they are washed away in the limit of long times. The presence of these non-classical correlations may be relevant to the engineering of entangled states for quantum computation. We highlight that these intriguing observations may depend on temperature, a question that we leave for future studies.
@@
V Assessment of the effective Hamiltonian method
We summarize the regime of applicability of the effective Hamiltonian theory. The effective Hamiltonian method prepares a system Hamiltonian that incorporates strong coupling effects through renormalization of energy parameters and the generation of bath-induced interaction terms. As such, by inspecting the effective Hamiltonian one can gain a direct understanding of expected strong coupling signatures. On the other hand, the coupling of the effective Hamiltonian to the residual bath is assumed weak and it is handled at the level of the Born-Markov approximation. Therefore, non-Markovian relaxation effects in the effective Hamiltonian dynamics are not taken into account. Based on this understanding, we summarize our assessment of the effective Hamiltonian theory:
(i) The theory properly captures the steady state behavior of spins in a boson bath, from the ultraweak to the ultrastrong coupling limits. We speculate that the method becomes more accurate when increasing the number of spins in the system at the level of equilibrium state description. It requires further assessment to understand the validity of the effective Hamiltonian method as the thermodynamic limit $N\to\infty$ is approached.